# Talk Isn’t Always Cheap: Understanding Failure Modes in Multi-Agent Debate

### v1

@@ Abstract
While multi-agent debate has been proposed as a promising strategy for improving AI reasoning ability, we find that debate can sometimes be harmful rather than helpful. The prior work has exclusively focused on debates within homogeneous groups of agents, whereas we explore how diversity in model capabilities influences the dynamics and outcomes of multi-agent interactions. Through a series of experiments, we demonstrate that debate can lead to a decrease in accuracy over time — even in settings where stronger (i.e., more capable) models outnumber their weaker counterparts. Our analysis reveals that models frequently shift from correct to incorrect answers in response to peer reasoning, favoring agreement over challenging flawed reasoning. These results highlight important failure modes in the exchange of reasons during multi-agent debate, suggesting that naive applications of debate may cause performance degradation when agents are neither incentivized nor adequately equipped to resist persuasive but incorrect reasoning.
Machine Learning, Multi-Agent Debate
@@ 1 Introduction
Large Language Model (LLM) agents have demonstrated remarkable problem-solving abilities across a wide array of complex reasoning tasks (Brown et al., 2020). Recently, a new line of research on interactive reasoning among multiple LLMs through debate has promoted the multi-agent debate framework as a promising approach to enhancing the reasoning and decision-making capabilities of LLM agents (Du et al., 2023; Chan et al., 2023; Liang et al., 2023; Khan et al., 2024). Various forms of multi-agent debate have been shown to improve performance on multiple arithmetic and strategic reasoning benchmarks (Du et al., 2023; Subramaniam et al., 2025), produce more truthful answers and evaluations (Chan et al., 2023; Khan et al., 2024), and enhance tasks such as machine translation (Liang et al., 2023) and negotiation (Fu et al., 2023). The core concept of these studies is that by engaging LLM agents through structured argumentation or discourse, we can facilitate the exchange of reasoning among different agents and guide them toward more accurate answers. However, most of these techniques incur significant computational overhead, as they are inference or test-time methods that rely on the strong zero-shot or few-shot capabilities of LLMs. Intuitively, greater exchanges of reasoning should lead to better decisions—allowing multiple agents to challenge flawed reasoning, highlight overlooked details, and reduce individual biases. But is this always the case?
@@
In this work, we show that the benefits of multi-agent debate are not as universal as commonly assumed. Through a series of empirical studies, we show that multi-agent debate can sometimes degrade performance, leading to worse final answers than those generated by a single agent acting alone. These failures are not rare edge cases, but arise systematically in settings where agents amplify each other’s errors – agreeing reflexively rather than challenging flawed reasoning. These findings hold true even when there is variation in the abilities of the participating LLM agents. For instance, we discover that introducing a weak or less capable (lower-performing) LLM agent into a debate with a strong or more capable (higher-performing) agent can detrimentally affect the debate outcome, producing results worse than if the agents had not engaged in discussion. The presence of a weaker agent disrupts the performance of the stronger agent. Moreover, in certain cases, the longer a debate continues, the more performance can degrade. In other words: talk isn’t always cheap – and in some cases, it’s actively harmful.
@@
We present a systematic evaluation of multi-agent debate across multiple tasks, showing that debate can sometimes harm group performance, particularly with heterogeneous LLM agents engaged in debate. Our findings challenge the prevailing narrative that more discussion between agents is inherently beneficial. Instead, we uncover several key factors that mediate the success or failure of debate, including task type and complexity, agent diversity and capability. In doing so, we offer a nuanced view of when and why debate helps – and when it hurts. Together, these results suggest that while debate remains a promising tool for improving model reasoning, its success is far from guaranteed.
@@ 2 Related Work
Debate and multi-agent reasoning. Multi-agent debate was initially proposed as method for scalable oversight problem where a judge or verifier can interject and elicit hidden contradictions, using structured back-and-forth conversations (Irving et al., 2018; Khan et al., 2024; Michael et al., 2023; Kenton et al., 2024). More recently, another form of multi-agent debate, sometimes referred to as multi-agent deliberation, investigates at leveraging different LLM agents to surface better answers by having them exchange reasoning via iterative discussion (Du et al., 2023; Chan et al., 2023; Liang et al., 2023; Subramaniam et al., 2025). Most of these studies focus on a homogeneous setting where all LLM agents utilize the same underlying language model, finding that this approach enhances accuracy across various Question-Answering (QA) tasks.
@@
Estornell & Liu (2024) examine the theoretical properties and effects of opinion diversity within the debate framework, reporting a“tyranny of the majority” effect. They found that if the majority of agents provide the same answer—regardless of its correctness—minority agents tend to conform, creating an echo chamber effect. Their analysis was limited to a homogeneous setting, whereas we extend our analysis to heterogeneous settings. Our findings reveal that even when the majority consists of stronger models, introducing a weaker model can diminish overall performance, thereby complementing their results.
Some works explores debates between agents of diverse nature or abilities. For instance, Estornell et al. (2025); Subramaniam et al. (2025) propose training LLM agents to debate collaboratively with distinct roles (actors/generators and critics), demonstrating that this approach can surpass previous unsupervised debate setups in reasoning benchmarks. Finally, studies such as Amayuelas et al. (2024) investigate whether the collaborative nature holds if an explicit adversary is introduced into the debate process—where the adversary actively seeks to reduce performance.
@@
These works suggest that, when agents constructively challenge each other, answers can improve. However, other studies caution that debate can fail when agents emphasize persuasion over truth. Agarwal & Khanna (2025) introduce a single-round debate on factual questions (using TruthfulQA) where one agent states a true answer calmly and another delivers a confident, emotional false answer. They show the LLM judge often chooses the persuasive falsehood with high confidence, suggesting that a vivid but incorrect argument can override a correct one. These results highlight a risk: unless the judge is well-calibrated, debate may amplify bluster, mirroring human misinformation scenarios.
@@
Collaborative multi-agent frameworks. Beyond explicit debates, many recent systems assume collaboration among LLMs improves reasoning (Li et al., 2023; Tran et al., 2025). For instance, Wu et al. (2023) provides a general multi-agent conversation framework: developers can define many agents (assistant, user-proxy, tools, etc.) that autonomously chat to solve complex tasks. These role-based and decentralized systems often yield richer interactions (e.g. multi-turn planning) than a single LLM alone.
@@
Sequential revision: Frameworks that utilize interactive reasoning at inference time in a sequential manner are employed for either self-refinement (Madaan et al., 2023) or self-consistency (Wang et al., 2022). Self-refinement involves iteratively revising or adapting a model’s responses based on previous outputs, prompting the model to intentionally reflect on its existing responses and correct any mistakes (Kamoi et al., 2024). Whereas works based on self-consistency often explicitly run multiple reasoning paths in parallel. For example, Wang et al. (2023) sample numerous independent chain-of-thought answers and select the most common answer. This aggregation of diverse reasoning paths significantly enhances the accuracy of arithmetic and commonsense QA by minimizing uncertainty. He et al. (2025) extend this idea: they spawn multiple “reactive” and “reflection” agent pairs, each exploring a different reasoning path, and then use a separate summarizer to aggregate them. Likewise, Yang et al. (2025) build a decentralized multi-agent planner where each LLM agent maintains its own memory (a hierarchical knowledge graph) and communicates via structured prompts. They find that these collaborative agents reach goals with  60% fewer steps than a lone agent, underscoring the value of structured cooperation.
@@
Collectively, prior work shows multi-agent debate and collaboration can amplify reasoning abilities, but also highlights key failure modes: judges may be fooled by rhetoric and human-like dynamics can bias group outputs. Our work builds on these insights by examining exactly when and why debate among LLMs helps or hurts performance on a variety of tasks.
@@ 3 Setting: Multi-agent debate
Let $\mathcal{Q}$ denote the a dataset of questions related to a task, where each question $q\in\mathcal{Q}$ in the task is natural language text. The objective is to generate an answer $a\in\mathcal{A}$ for any given input question $q\in\mathcal{Q}$, where $\mathcal{A}$ is the set of possible answers for that task. We assume that there exists a ground truth answer $a^{\star}\in\mathcal{A}$ for each question $q\in\mathcal{Q}$ which is denoted by $f^{gt}:\mathcal{Q}\rightarrow\mathcal{A}$, i.e., $f^{gt}(q)=a^{\star}$.
Single-Agent Setting: In the single agent setting, each agent uses an undelrying LLM $l:\mathcal{Q}\rightarrow\mathcal{A}$ to generate answer $a\sim l(q)$ where $a\in\mathcal{A}$ denotes the answer generated by the LLM in response to the question $q$. The main metric of interest is accuracy which is calculated as $\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbbm{1}[l(q)=f^{gt}(q)]$.
@@
Multi-Agent Debate: We follow the multi-agent debate framewor from (Du et al., 2023; Subramaniam et al., 2025) that involves initially posing a question to a group of LLM agents. After the initial responses are generated, the debate process then iteratively revisits the question for each agent, contextualizing input through the collective responses from previous rounds. Specifically, each agent is generates a new response to the question based on its prior response and the summarized responses from the other agents. The final answer is based by majority vote among all the agents.
Formally, we have a group of $N$ agents, each with their own LLM $l_{i}$ that are presented with a question $q\in\mathcal{Q}$ and tasked with generating an answer $a\in\mathcal{A}$. We use $d(l_{1},\ldots,l_{n}):\mathcal{Q}\rightarrow\mathcal{A}$ to denote the debate procedure that takes input some question and set of LLM agents and generates an answer. The debate procedure runs over mutliple rounds, where at any given round $t$, each agent $i$ uses their underlying LLM $l_{i}$ to iteratively generate an answer in the following manner:
@@
    Starting Round ($t=1$): Each agent $i$ generates an response $g_{i}\in\mathcal{G}$ via $l_{i}$ with an optional additional prompt, $S:\mathcal{Q}\times\mathcal{P}_{\texttt{starting}}\rightarrow\mathcal{A}$, or $g_{i}^{1}\sim l_{i}(\mathcal{P}_{\texttt{starting}}(q))$. Here $\mathcal{G}$ denotes the set of possible generations from LLM.
@@
    Debate Rounds ($t=2,\ldots,T$): For the subsequent rounds, each agent $i$ is given the question $q$ and responses from the other agents from the previous round. Let $o_{i}^{t}=\left\{g_{j}^{t-1}\right\}_{j\neq i}$ denote the outputs from the other agents from previous round, then the goal of agent $i$ for this round is to generate an updated response $g^{t}_{i}\in\mathcal{G}$ that takes into account the responses from the other agents and it’s own response from the previous round $g^{t-1}_{i}$, i.e.,
@@
This procedure runs over for $T$ rounds, where the final output of the debate is the set of responses $\{g^{T}_{i}\}_{i=1}^{n}$. Then the majority response across all agents is selected the final answer. Let $\texttt{majority}:\mathcal{G}^{N}\rightarrow\mathcal{A}$ denote the majority voting and filtering for responses across all agents, then the final answer is given by $a=\texttt{majority}(\{g^{T}_{i}\}_{i=1}^{n})$.
@@ 4 Experimental Setup
We describe the set of tasks, models and prompts used in our experiments in this section.
@@ 4.1 Datasets
CommonSenseQA: The CommonSenseQA dataset (Talmor et al., 2019) consist of multiple-choice questions with complex semantics that often require prior knowledge to answer correctly. The dataset is intended to test for prior common-sense knowledge encoded within LLMs and check for common misconceptions.
@@ 4.2 Models
We used three models from distinct model families in our experiments: GPT-4o-mini (OpenAI, 2024), LLaMA-3.1-8B-Instruct (et al., 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023). To align with prior work on multi-agent debate (Du et al., 2023; Subramaniam et al., 2025), we ran all experiments and models with default temperature parameter, the $\texttt{top\_p}=0.9$, maximum generation length of $2048$ tokens and $T=2$ rounds of debate. We take 100 random samples for each task from the dataset and report result over 5 random seeds.

### v2

@@ Abstract
While multi-agent debate has been proposed as a promising strategy for improving AI reasoning ability, we find that debate can sometimes be harmful rather than helpful. Prior work has primarily focused on debates within homogeneous groups of agents, whereas we explore how diversity in model capabilities influences the dynamics and outcomes of multi-agent interactions. Through a series of experiments, we demonstrate that debate can lead to a decrease in accuracy over time — even in settings where stronger (i.e., more capable) models outnumber their weaker counterparts. Our analysis reveals that models frequently shift from correct to incorrect answers in response to peer reasoning, favoring agreement over challenging flawed reasoning. We perform additional experiments investigating various potential contributing factors to these harmful shifts – including sycophancy, social conformity, and model and task type. These results highlight important failure modes in the exchange of reasons during multi-agent debate, suggesting that naive applications of debate may cause performance degradation when agents are neither incentivised nor adequately equipped to resist persuasive but incorrect reasoning.
@@ 1 Introduction
Large Language Model (LLM) agents have demonstrated remarkable problem-solving abilities across a wide array of complex reasoning tasks (Brown et al., 2020). Recently, a new line of research on interactive reasoning among multiple LLMs through debate has promoted the multi-agent debate framework as a promising approach to enhancing the reasoning and decision-making capabilities of LLM agents (Du et al., 2023; Chan et al., 2023; Liang et al., 2023; Khan et al., 2024). Various forms of multi-agent debate have been shown to improve performance on multiple arithmetic and strategic reasoning benchmarks (Du et al., 2023; Subramaniam et al., 2025), produce more truthful answers and evaluations (Chan et al., 2023; Khan et al., 2024), and enhance tasks such as machine translation (Liang et al., 2023) and negotiation (Fu et al., 2023). The core concept of these studies is that by engaging LLM agents through structured argumentation or discourse, we can facilitate the exchange of reasoning among different agents and guide them toward more accurate answers. Intuitively, greater exchanges of reasoning should lead to better decisions—allowing multiple agents to challenge flawed reasoning, highlight overlooked details, and reduce individual biases. But is this always the case?
@@
In this work, we show that the benefits of multi-agent debate are not as universal as commonly assumed. Through a series of empirical studies, we show that multi-agent debate can sometimes degrade performance, leading to worse final answers than those generated by a single agent acting alone. These failures are not rare edge cases, but arise systematically in settings where agents amplify each other’s errors – agreeing reflexively rather than challenging flawed reasoning. These findings hold even when there is variation in the abilities of the participating LLM agents. For instance, we discover that introducing a weak or less capable (lower-performing) LLM agent into a debate with a strong or more capable (higher-performing) agent can detrimentally affect the debate outcome, producing results worse than if the agents had not engaged in discussion. The presence of a weaker agent disrupts the performance of the stronger agent. Moreover, in certain cases, the longer a debate continues, the more performance can degrade. In other words: talk isn’t always cheap – and in some cases, it’s actively harmful.
@@
We present a systematic evaluation of multi-agent debate across multiple tasks, showing that debate can sometimes harm group performance, particularly with heterogeneous LLM agents engaged in debate. Our findings challenge the prevailing narrative that more discussion between agents is inherently beneficial. Instead, we uncover several key factors that mediate the success or failure of debate, including task type and complexity, agent diversity and capability, and social influence. In doing so, we offer a nuanced view of when and why debate helps – and when it hurts. Together, these results suggest that while debate remains a promising tool for improving model reasoning, it should be applied carefully to ensure safety on the task and setting of interest.
Statement of Contributions. We summarize our contributions as follows:
  1. 1.
    We conduct a comprehensive evaluation of the effectiveness of the multi-agent debate framework across three different datasets. In Section˜5.1, we demonstrate that debate may degrade performance compared to majority voting. Additionally, in Section˜5.2 we show that the performance may progressively deteriorate as the debate progresses.
  2. 2.
    We extend the debate framework beyond the homogeneous setting to investigate the effect of diverse agent populations. Contrary to the prevalent belief that LLM agents are inherently collaborative – suggesting that a mixture of diverse models improves response quality when they can access outputs from other models, even if those outputs are of lower quality (Wang et al., 2024) – we find this assumption does not hold in multi-agent debates. In Section˜5, we observe that the presence of weaker agents can negatively affect performance.
  3. 3.
    Our analysis in Section˜6 reveals that a significant portion of correct answers become corrupted during debate. We investigate this phenomenon from various perspectives, such as the effect of sequential revision (Section˜6.1), social influence (Section˜6.2), and sycophancy (Section˜6.3). These insights opens avenues for future research focused on enhancing reasoning exchange in multi-agent systems.
@@ 2 Related Work
Debate and multi-agent reasoning. Multi-agent debate was initially proposed as method for the scalable oversight problem where a judge or verifier can interject and elicit hidden contradictions, using structured back-and-forth conversations (Irving et al., 2018; Khan et al., 2024; Michael et al., 2023; Kenton et al., 2024). More recently, another form of multi-agent debate, sometimes referred to as multi-agent deliberation, investigates leveraging different LLM agents to surface better answers by having them exchange reasoning via iterative discussion (Chan et al., 2023; Liang et al., 2023; Subramaniam et al., 2025). Most of these studies focus on a homogeneous setting where all LLM agents utilize the same underlying language model (Du et al., 2023) or a model of similar ability (Yao et al., 2025), finding that this approach enhances accuracy across various Question-Answering (QA) tasks.
@@
Estornell and Liu (2024) examine the theoretical properties and effects of opinion diversity within the debate framework, reporting a “tyranny of the majority” effect. They found that if the majority of agents provide the same answer – regardless of its correctness – minority agents tend to conform, creating an echo chamber effect. Additionally, they propose a theoretical result indicating that diversity of model abilities should improve overall debate performance; we show empirical evidence that this is often not true in practice.
@@
Some works explore debates between agents of diverse nature or abilities. For instance, Estornell et al. (2025); Subramaniam et al. (2025) propose training LLM agents to debate collaboratively with distinct roles (actors/generators and critics), demonstrating that this approach can surpass previous unsupervised debate setups in reasoning benchmarks. Finally, studies such as Amayuelas et al. (2024) investigate whether the collaborative nature holds if an explicit adversary is introduced into the debate process—where the adversary actively seeks to reduce performance. These works suggest that, when agents constructively challenge each other, answers can improve. However, other studies caution that debate can fail when agents emphasize persuasion over truth. Agarwal and Khanna (2025) introduce a single-round debate on factual questions (using TruthfulQA) where one agent states a true answer calmly and another delivers a confident, emotional false answer. They show the LLM judge often chooses the persuasive falsehood with high confidence, suggesting that a vivid but incorrect argument can override a correct one. These results highlight a risk: unless the judge is well-calibrated, debate may amplify bluster, mirroring human misinformation scenarios. Yao et al. (2025) focus on investigating sycophancy in the debate setting, showing that agent disagreement rate decreases as debate progresses, and that this observation is correlated with performance degradation.
@@
Collaborative multi-agent frameworks. Beyond explicit debates, many recent systems assume collaboration among LLMs improves reasoning (Li et al., 2023; Wang et al., 2024; Tran et al., 2025). For instance, Wu et al. (2023) provides a general multi-agent conversation framework: developers can define many agents (assistant, user-proxy, tools, etc.) that autonomously chat to solve complex tasks. These role-based and decentralized systems often yield richer interactions (e.g. multi-turn planning) than a single LLM alone.
@@
Sequential revision. Frameworks that utilize interactive reasoning at inference time in a sequential manner are employed for either self-refinement (Madaan et al., 2023) or self-consistency (Wang et al., 2022). Self-refinement involves iteratively revising or adapting a model’s responses based on previous outputs, prompting the model to intentionally reflect on its existing responses and correct any mistakes (Kamoi et al., 2024). Works based on self-consistency often explicitly run multiple reasoning paths in parallel. For example, Wang et al. (2023) sample numerous independent chain-of-thought answers and select the most common answer. This aggregation of diverse reasoning paths significantly enhances the accuracy of arithmetic and commonsense QA by minimizing uncertainty. He et al. (2025) extend this idea: they spawn multiple “reactive” and “reflection” agent pairs, each exploring a different reasoning path, and then use a separate summarizer to aggregate them. Likewise, Yang et al. (2025) build a decentralized multi-agent planner where each LLM agent maintains its own memory (a hierarchical knowledge graph) and communicates via structured prompts. They find that these collaborative agents reach goals with  60% fewer steps than a lone agent, underscoring the value of structured cooperation.
@@
Collectively, prior work shows that multi-agent debate and collaboration can amplify reasoning abilities, but also highlights key failure modes: judges may be fooled by rhetoric, and human-like dynamics can bias group outputs. Our work builds on these insights by expanding the analysis to heterogeneous debate settings, then performing a deep dive into the potential factors affecting when debate helps – or hurts – LLM performance on a variety of tasks.
@@ 3 Setting: Multi-agent debate
Let $\mathcal{Q}$ denote a dataset of questions related to a task, where each question $q\in\mathcal{Q}$ in the task is natural language text. The objective is to generate an answer $a\in\mathcal{A}$ for any given input question $q\in\mathcal{Q}$, where $\mathcal{A}$ is the set of possible answers for that task. We assume that there exists a ground truth answer $a^{\star}\in\mathcal{A}$ for each question $q\in\mathcal{Q}$ which is denoted by $f^{gt}:\mathcal{Q}\rightarrow\mathcal{A}$, i.e., $f^{gt}(q)=a^{\star}$.
Single-Agent Setting: In the single agent setting, each agent uses an underlying LLM $l:\mathcal{Q}\rightarrow\mathcal{A}$ to generate answer $a\sim l(q)$ where $a\in\mathcal{A}$ denotes the answer generated by the LLM in response to the question $q$. The main metric of interest is accuracy which is calculated as $\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathbbm{1}[l(q)=f^{gt}(q)]$.
@@
Multi-Agent Debate: We follow the multi-agent debate framework from Du et al. (2023); Subramaniam et al. (2025) which involves initially posing a question to a group of LLM agents. After the initial responses, which include answers plus reasoning, are generated, the debate process then iteratively revisits the question for each agent, contextualizing input through the collective responses from previous rounds. Specifically, each agent generates a new response to the question based on its prior response and the summarized responses from the other agents. The final answer is based on the majority vote among all agents.
Formally, we have a group of $N$ agents, each with their own LLM $l_{i}$, that are presented with a question $q\in\mathcal{Q}$ and tasked with generating an answer $a\in\mathcal{A}$. We use $d(l_{1},\ldots,l_{n}):\mathcal{Q}\rightarrow\mathcal{A}$ to denote the debate procedure that takes as input a question and a set of LLM agents and generates an answer. The debate procedure runs over multiple rounds, where in any given round $t$, each agent $i$ uses their underlying LLM $l_{i}$ to iteratively generate an answer in the following manner:
@@
    Starting Round ($t=0$): Each agent $i$ generates a response $g_{i}\in\mathcal{G}$ via $l_{i}$ using a starting prompt $\mathcal{P}_{\texttt{starting}}$, i.e. $g_{i}^{0}\sim l_{i}(\mathcal{P}_{\texttt{starting}}(q))$. Here $\mathcal{G}$ denotes the set of possible generations from LLM.
@@
    Debate Rounds ($t=1,\ldots,T$): For the subsequent rounds, each agent $i$ is given the question $q$ and responses from the other agents from the previous round. Let $o_{i}^{t}=\left\{g_{j}^{t-1}\right\}_{j\neq i}$ denote the outputs from the other agents from previous round. Then the goal of agent $i$ for this round is to generate an updated response $g^{t}_{i}\in\mathcal{G}$ that takes into account the responses from the other agents and its own response from the previous round $g^{t-1}_{i}$, i.e.,
@@
This procedure runs for $T$ rounds, where the final output of the debate is the set of responses $\{g^{T}_{i}\}_{i=1}^{n}$. Then the majority response across all agents is selected as the final answer. Let $\texttt{majority}:\mathcal{G}^{N}\rightarrow\mathcal{A}$ denote the majority voting and filtering for responses across all agents, then the final answer is given by $a=\texttt{majority}(\{g^{T}_{i}\}_{i=1}^{n})$. Accuracy is evaluated with respect to the final for the group of debating agents.
@@ 4 Experimental Setup
In this section, we describe the set of tasks, models, and prompts used in our experiments.
@@ 4.1 Datasets
CommonSenseQA (CSQA): The CommonSenseQA dataset (Talmor et al., 2019) consists of multiple-choice questions with complex semantics that often require prior knowledge to answer correctly. The dataset is intended to test for prior common-sense knowledge encoded within LLMs and checks for common misconceptions. Additionally, this task (common-sense reasoning) was not evaluated in prior work on multi-agent debate, making it a good testbed for evaluating generalization of these debate approaches to different tasks.
@@ 4.2 Models
We used three models from distinct model families in our experiments: GPT-4o-mini (OpenAI, 2024), LLaMA-3.1-8B-Instruct (et al., 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023). To align with prior work on multi-agent debate (Du et al., 2023; Subramaniam et al., 2025), we ran all experiments and models with the default temperature parameter, the $\texttt{top\_p}=0.9$, maximum generation length of $2048$ tokens, and $T=2$ rounds of debate. We take 100 random samples for each task from the dataset and report the result over 5 random seeds.